OpenStudy (anonymous):

How many numbers are twice as far from 5 as they are from 8 and what are they?

OpenStudy (valpey):

when we say "far" we usually mean something like the absolute value of difference. So let one of these numbers be called "n". Then we'd have: 2|n-5| = |n-8|

OpenStudy (valpey):

Now a cool thing about absolute values is that when you square them you can erase the absolute value function and manipulate them with algebra: \[(2|n-5|)^2 = (|n-8|)^2\]\[2^2(n-5)^2=(n-8)^2\]

OpenStudy (valpey):

Now try to solve for n.

OpenStudy (anonymous):

i tried that but i got \[3n ^{2}= 39\] which eventually got me to about 3.6

OpenStudy (anonymous):

it just doesnt seem like the right answer

OpenStudy (anonymous):

Pepper, you seem to be doing your algebra incorrectly...it should look something like this (starting from Valpey's equation): \[\begin{align} &2^2(n-5)^2=(n-8)^2\\ &4(n^2-10n+25)=n^2-16n+64\\ &4n^2-40n+100=n^2-16n+64\\ &3n^2-24n+36=0\\ &(3n-18)(n-2)=0\\ &n=6,\;2 \end{align}\]

OpenStudy (valpey):

Oops..."twice as far from five..." It sounds like it must be: \[2^2(n-8)^2=(n-5)^2\]instead.

OpenStudy (anonymous):

Yep, you're right...this is what I get for not reading the problem.

OpenStudy (anonymous):

thanks i just realized that i didnt multiply right stupidd...

OpenStudy (anonymous):

wait is n the number of numbers that are twice as far from 5 as they are from 8?

OpenStudy (anonymous):

do you guys know if the answer is 7?